OpenStudy (trojanpoem):

If the tangent line to the curve : y = x^3 - 3x^2 makes an obtuse angle with the positive x-axis. then :

OpenStudy (trojanpoem):

\[x \in ....\]

OpenStudy (anonymous):

what type of answer do you want not direct? not steps?

OpenStudy (anonymous):

but you said not steps was it a typo?

OpenStudy (trojanpoem):

a) [0,2] b) ]0,2[ c) Z- [0,2] d) Z - ]0,2[ Looking for steps.

OpenStudy (anonymous):

the question itself is sort of ends with then? what

OpenStudy (trojanpoem):

If the tangent line to the curve : y = x^3 - 3x^2 makes an obtuse angle with the positive x - axis then x∈.... a) [0,2] b) ]0,2[ c) Z- [0,2] d) Z - ]0,2[

OpenStudy (anonymous):

well looking at the plot (0,1) would be my answer

OpenStudy (anonymous):

i plotted the curve and looked at the tangent line ---wait for another reply maybe they will see it different

OpenStudy (irishboy123):

i think you need to plot the curve, and then draw the tangent line in a few times. you will see when it makes an obtuse angle, as opposed to an acute angle, with the x direction....

OpenStudy (trojanpoem):

Isn't there any solution without plotting it and even if so plot it and tell me which answer you think is right ( I got the final answer ).

OpenStudy (irishboy123):

the tangent line only points downwards between 0 < x < 2. at all other times it points up (makes an acute angle with x axis). https://www.wolframalpha.com/input/?i=plot+y+%3D+x%5E3+-+3+x%5E2

OpenStudy (trojanpoem):

Right the answer is ]0, 2[ but I want a solution without plotting it.

OpenStudy (trojanpoem):

I got idea, how about using maxima and mimima and check where it's decreasing.

OpenStudy (trojanpoem):

dy/dx = 3x^2 - 6x x = 0 x= 2 in crease Z- [0,2] decrease ]0,2[ I got it.

OpenStudy (trojanpoem):

Do you agree with this solution ?

OpenStudy (phi):

Just to be sure, I would sketch the graph. but "obtuse angle" means negative slope. so solve 3x^2 - 6x < 0 3x(x-2) < 0 we know x >=0 so 3x is positive, i.e. x>0 (x =0 gives zero slope, which means the line does not intersect the x-axis) thus to get a negative product we require x-2 < 2 x <2 and 0 < x < 2